OpenStudy (anonymous):

A fruit company delivers its fruit in two types of boxes, large and small. A delivery of 4 large boxes and 8 small boxes has a total weight of 105 kilograms. A delivery of 2 large boxes and 3 small boxes has a total weight of 46 kilograms.How much does each type of box weigh?

OpenStudy (anonymous):

4L+8S=105 and 2L+3S=46 now solve the system of linear equations to find the values of L and S.

OpenStudy (anonymous):

How do I do that @hullsnipe

OpenStudy (anonymous):

Solve one of the equations for "L" then substitute that value into the other equation and solve for "S", you will get a number. Now plug that number in for S into either equation and solve for the number of L. The first step I mentioned is as follows: \[4L+8S=105\] \[4L = 105-8S\] \[L=\frac{ 105-8S }{4 }\] Now use this to continue on.